OpenStudy (anonymous):

what value of k solves the equation k^-5 = 1/32

OpenStudy (anonymous):

\[k^{-5}=\frac{1}{k^5}=\frac{1}{32}\] which tells you \[k^5=32\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

What value of m solves the equation 2^m = 1/8

OpenStudy (anonymous):

1/32^-(1/5) Or in other words to -5 root of 1/32 Alternatively the 5 root of 32 in other words 32^(1/5)

OpenStudy (anonymous):

x^m = y Log base x of y = m

OpenStudy (anonymous):

the answer is 2. i just took the test, trust me.

jabez177 (jabez177):

Thank you @amberosales!

OpenStudy (anonymous):

Lol no prob @jabez177